On heating to dryness, 3.10g of a hydrated salt produced 2.02g of anhydrous salt.The formula of the hydrated salt is represented as M.xH2O. If the molecular mass of salt M is 202g/mol , what is the value of x ?
ans: A. 4
B. 5
C. 6 ( )
D. 7
1
Expert's answer
2015-11-23T10:34:49-0500
Solution
M*xH2O = M + xH2O
n(H2O) = (3.10-2.02)/18 = 0.06
n(M*xH2O) = 3.10/(202 + 18x)
n(M*xH2O) = x*n(H2O)
3.10/(202+18x) = x*0.06
0.015 + 0.17x = 0.06x
X = 7
Answer
D. 7